# When Machine Unlearning Meets Retrieval-Augmented Generation (RAG): Keep Secret or Forget Knowledge?

### v1

@@ Abstract
The deployment of large language models (LLMs) like ChatGPT and Gemini has shown their powerful natural language generation capabilities. However, these models can inadvertently learn and retain sensitive information and harmful content during training, raising significant ethical and legal concerns. To address these issues, machine unlearning has been introduced as a potential solution. While existing unlearning methods take into account the specific characteristics of LLMs, they often suffer from high computational demands, limited applicability, or the risk of catastrophic forgetting. To address these limitations, we propose a lightweight unlearning framework based on Retrieval-Augmented Generation (RAG) technology. By modifying the external knowledge base of RAG, we simulate the effects of forgetting without directly interacting with the unlearned LLM. We approach the construction of unlearned knowledge as a constrained optimization problem, deriving two key components that underpin the effectiveness of RAG-based unlearning. This RAG-based approach is particularly effective for closed-source LLMs, where existing unlearning methods often fail. We evaluate our framework through extensive experiments on both open-source and closed-source models, including ChatGPT, Gemini, Llama-2-7b-chat-hf, and PaLM 2. The results demonstrate that our approach meets five key unlearning criteria: effectiveness, universality, harmlessness, simplicity, and robustness. Meanwhile, this approach can extend to multimodal large language models and LLM-based agents.
@@ I Introduction
Large language models (LLMs) have transformed the field of natural language generation, playing a vital role in advancing technology across various industries [1, 2]. Models like the GPT series, which are large-scale neural networks, have learned the complexities of language—both semantics and syntax—by training on vast datasets, enabling them to excel in a wide range of language tasks [3, 4]. Currently, the training of LLMs typically follows a ‘pre-training and fine-tuning’ paradigm [5, 6]. However, both stages may inadvertently result in the model learning sensitive information or harmful content [7, 8, 9], which raises concerns about privacy violations, copyright infringement, or the generation of harmful content.
@@
To address the privacy or harmful content concerns, machine unlearning has been proposed as a method to remove the influence of specific data from a model without requiring full retraining [13]. Unlike traditional unlearning in classification models, fully eliminating the contribution of target knowledge in LLM faces two challenges. The key challenge is the difficulty of modifying the parameters in LLMs due to the huge volumes of parameters. Second, the output space of LLMs is much larger than the classification space, which hinders the verification of LLM unlearning. The two challenges block the implementation of exact LLM unlearning.
@@
Therefore, approximate unlearning for LLM has become a research hotspot [11, 12, 14]. It hypothesizes that the unlearned LLM differs from the regular model in the target knowledge and behaves similarly in the remaining knowledge. Although model owners can not remove the target knowledge in LLM, they can hide it or map it to irrelevant content, such as poisoning the target, shifting the logits and modifying the input. However, these LLM unlearning schemes have inherent weaknesses. First, the schemes involving fine-tuning LLM introduce considerable computational overhead, like gradient ascent, so they cannot adapt to frequent unlearning requests. Second, closed-source LLMs only provide their interfaces, making some schemes unsuitable. Third, LLMs possess emergent abilities, which means that forgetting must extend beyond specific data to include related content. Shumailov et al. [15] discovered that LLMs can recall previously forgotten knowledge through their in-context abilities, exposing a vulnerability in existing unlearning schemes to this ‘un-unlearning’ phenomenon. Furthermore, some unlearning schemes may reduce the overall utility of the model, potentially leading to catastrophic forgetting [16]. In Table I, we list three representative LLM unlearning schemes and their weaknesses.
@@
To tackle above weaknesses of LLM unlearning, we find that Retrieval-Augmented Generation (RAG) is a new emerged technology that does not need to change any parameter but with adjusting the output of LLMs [17]. This special characteristic can possibly be applied in the machine unlearning. In general, the RAG workflow comprises two primary components: an information retrieval module and a text generation module [18]. The retrieval module first retrieves relevant information from a knowledge base based on the input. The LLM then uses this retrieved information to generate the final response. This process can effectively improve the quality of outputs, reduce the likelihood of misinformation and hallucinations, and, most importantly, hide the target information to implement the unlearning without revising the parameters.
@@
As illustrated in Figure 1 (a), RAG can revise the misinformation knowledge in the external knowledge base, effectively correcting it in LLM. When regarding misinformation as a forgotten target, the scalability of RAG technology offers a promising solution for LLM unlearning. In Figure 1 (b), through adding confidentiality requirements of the target knowledge in the knowledge base, RAG can hide the forgotten target’s information without requiring access to the unlearned LLM. Figure 1 (c) indicates that an ideal unlearning scheme would remove the target knowledge in LLM, thus not generating its related content. Fortunately, leveraging RAG to keep the forgotten target confidential can function similarly to unlearning it.
@@
Combined with the idea of modifying the input, we take advantage of RAG technology to propose a novel approximate LLM unlearning framework. Specifically, it treats unlearning the target data as a constrained optimization problem for constructing its unlearned knowledge. The knowledge consists of two main components: a retrieval component and a constraint component. For each target, we use a heuristic method to build relevant knowledge (i.e., retrieval component) and confidentiality requirements (i.e., constraint component). To improve the effectiveness of unlearning, we apply prompt engineering to optimize both components. Our approach only modifies the knowledge base to implement unlearning. This flexibility allows our framework to be applied to both open-source and closed-source models, thereby adapting to closed-source models and overcoming high high computational demand. Furthermore, we can utilize the unlearned LLM to acquire knowledge related to the target data, thereby keeping the extended knowledge confidential. This can address the ‘un-unlearning’ phenomenon. Meanwhile, RAG is like an external plug-in that will not affect the model utility when inputs are not related to the knowledge base. This ensures that our scheme can easily land in various LLMs with minimal side effects.
In our experiment, we focused on two key unlearning objectives: trained samples and learned concepts. RAG-based unlearning was implemented in both open-source and closed-source models, including GPT-4o, Gemini, and Llama-2-7b-chat-hf. Additionally, We evaluated our unlearning framework comprehensively across five key dimensions: effectiveness, universality, harmlessness, simplicity, and robustness, and found that the framework performed well in each dimension. Moreover, we explored the resistance of our scheme against jailbreak attacks and prompt injection attacks to show that our method can efficiently unlearn the target samples. Furthermore, we extended RAG-based unlearning to multimodal large language models (MLLMs) and LLM-based agents, and achieved effective forgetting effects.
In summary, RAG-based unlearning enables model owners to control the knowledge boundaries of LLMs by managing external knowledge bases, thereby safeguarding privacy, protecting copyrighted data, and eliminating harmful content. Our primary contributions are as follows:
  • We are the first to propose an end-to-end unlearning framework leveraging RAG technology, which can be applied to many LLMs.
@@ II-A1 The Definition of Machine Unlearning
There are two types of machine unlearning. The first is exact unlearning, and it guarantees that the distribution of $F_{U}$ and that of a model retrained on $D_{R}$ are indistinguishable. In practice, exact unlearning can only be implemented for tiny-scale models. Therefore, approximate unlearning comes up and only guarantees that the activation results of $F_{U}$ and that of a retrained model are indistinguishable. Clearly, LLMs are suitable for approximate unlearning due to their complex architectures.
Notably, the output of LLMs differs from that of traditional models, Definition 1 cannot be extended to LLM unlearning. We will detail it in Section III-B.
@@ II-B1 Machine Unlearning
Machine unlearning has become a crucial technique for addressing various safety issues, such as privacy protection and copyright protection. It originates from the right to be forgotten, which grants users the right to request that a model provider remove the influence of its data from the model. Currently, machine unlearning has expanded into various fields, including image classification [20, 21], recommendation systems [22], and graph neural networks [23]. In small-scale models, unlearning requests include sample unlearning [24], class unlearning [25], and feature unlearning [26]. To achieve these goals, data reorganization and model manipulation are employed. In the first approach, model providers will delete or modify the target data. For example, Bourtoule et al. developed a machine unlearning framework called SISA [20]. It works by partitioning the training data into several non-overlapping subsets (i.e., shards) and independently training models on each shard. When an unlearning request is received, SISA simply retrains the shard containing the target data. For model manipulation, model shifting and model pruning are used to affect the knowledge boundary regarding the target data. For instance, Wang et al. [27] proposed a class-unlearning scheme by quantifying the class discrimination of different channels in convolutional neural networks.
@@ II-B2 Existing LLM Unlearning
Different from machine unlearning, LLM unlearning remains in an exploratory phase. There are three types of LLM unlearning schemes: retraining-based, offset-based and input-based. In the first manner, Yao et al. [10] used a gradient ascent algorithm to increase the model’s loss on the target, causing the output to deviate from the original content. Similarly, Eldan et al. [14] re-labeled the forgotten data as non-sensitive responses, thereby minimizing their likelihoods. These unlearning schemes require accessing and manipulating LLMs and cannot be applied to closed-source LLMs. In the second manner, Huang et al. [11] employed logit shifting with two small-scale models to modify the LLM’s output. When given a forgotten sample, the shift prevents the LLM from generating sensitive information. However, this scheme relies on the performance of small models and requires precise logit shifting calculations. Input-based unlearning is popular, Pawelczyk et al. [12] leveraged the contextual capabilities of LLMs to implement an in-context unlearning method. They guided the model’s behavior by constructing counterfactual instances of the target data. However, this scheme increases the risk of privacy leaks, as it requires storing sensitive instances. Further, the LLM can not understand the counterfactual instances in most cases, hindering in-context unlearning.
@@ II-B3 Verification for Machine Unlearning
The purpose of verification is to ensure that the model cannot reveal information about the target data after unlearning operations. The model should closely resemble its counterpart that has never seen the target data. Currently, several verification methods for machine unlearning exist, including retraining time [20], privacy and security attacks [28, 29], and accuracy [30]. The first manner measures how quickly the unlearned model can relearn the forgotten data. If the model can recover its original performance in a very short time, this suggests that it may still retain information about the forgotten data. Furthermore, Graves et al. [28] employed membership inference attacks (MIAs) to verify machine unlearning. In a successful demonstration of unlearning, MIA determines the target data to be ‘in’ for the pre-unlearned model and ‘out’ for the post-unlearned model. Similarly, Guo et al. [29] used backdoored data as the target and calculated the drop of attack success rate. Literature [30] used an accuracy-based verification method, comparing the output of the model after unlearning with that of a model that has never seen the target data. Although these verification methods are intuitive, they cannot guarantee that the model definitely forgets the target data in all scenarios, particularly in LLMs. We will detail the verification for LLM unlearning in SectionV-A.
@@ II-B4 Retrieval-Augmented Generation
RAG in LLMs is an enhancement method that combines retrieval and generation techniques. Compared to regular LLMs, RAG-based models are particularly effective at handling complex tasks, especially in contexts that require detailed background knowledge and specific facts. Specifically, by introducing a retrieval process, RAG-based models can incorporate the most relevant information related to input prompts, reducing factual errors during the generation process. Additionally, RAG techniques are well-suited for handling multi-turn dialogues [31]. Currently, some researchers improved RAG techniques regarding retrieval process and generation process. Jiang et al. [32] optimized the retrieval algorithm for faster and more accurate information extraction. While Siriwardhana et al. [33] focused on the generation process to better integrate and utilize the retrieved knowledge.
@@ III Problem Definition
$BK$ The benign knowledge base
$UK$ The unlearned knowledge base
@@ III-A Notations
III-B The Definition of LLM Unlearning
According to recent works [12, 10], LLM unlearning has two key objectives: sample unlearning and concept unlearning.
@@ III-B The Definition of LLM Unlearning
Theoretically, an ideal unlearning scheme would retrain $A(;\theta)$ on the training corpus, excluding Esample or Econcept. However, this naive scheme is impractical due to its complexity. In this work, we adopt RAG technology to implement an approximation of LLM unlearning.
@@ III-C Threat Model
In contrast to traditional unlearning, we consider real-world unlearning scenarios for LLMs, including privacy protection, copyright protection, and removing harmful output. The first scenario primarily involves sample unlearning, while the latter two involve concept unlearning. As shown in Figure 3, both types of unlearning requests typically occur after the model deployment. Specifically, a model provider trains an LLM and releases it. In this case, a data contributor may request to withdraw its samples (i.e., partial training data) or concepts (a set of relevant training data). The model provider then uses LLM unlearning schemes to remove the influence of the target data on the model. After unlearning, the data contributor needs to verify whether the unlearned model has forgotten the target.
Under the threat model, our scheme requires only a model maintainer or a third party with minimal capability to perform the unlearning process. In the sample unlearning scenario, the model maintainer can access the trained samples that need to be forgotten and modify the input prompts. However, this process does not require access to the parameters of LLM. In the concept unlearning scenario, the model maintainer has similar capabilities as described above but only has knowledge of the forgotten concepts, rather than the related samples.

### v2

@@ Abstract
The deployment of large language models (LLMs) like ChatGPT and Gemini has shown their powerful natural language generation capabilities. However, these models can inadvertently learn and retain sensitive information and harmful content during training, raising significant ethical and legal concerns. To address these issues, machine unlearning has been introduced as a potential solution. While existing unlearning methods take into account the specific characteristics of LLMs, they often suffer from high computational demands, limited applicability, or the risk of catastrophic forgetting. To address these limitations, we propose a lightweight behavioral unlearning framework based on Retrieval-Augmented Generation (RAG) technology. By modifying the external knowledge base of RAG, we simulate the effects of forgetting without directly interacting with the unlearned LLM. We approach the construction of unlearned knowledge as a constrained optimization problem, deriving two key components that underpin the effectiveness of RAG-based unlearning. This RAG-based approach is particularly effective for closed-source LLMs, where existing unlearning methods often fail. We evaluate our framework through extensive experiments on both open-source and closed-source models, including ChatGPT, Gemini, Llama-2-7b-chat, and PaLM 2. The results demonstrate that our approach meets five key unlearning criteria: effectiveness, universality, harmlessness, simplicity, and robustness. Meanwhile, this approach can extend to multimodal large language models and LLM-based agents.
@@ I Introduction
Large language models (LLMs) have transformed the field of natural language generation, playing a vital role in advancing technology across various industries [1, 2]. Models like the GPT series, which are large-scale neural networks, have learned the complexities of language—both semantics and syntax—by training on vast datasets, enabling them to excel in a wide range of language tasks [3, 4]. Currently, the training of LLMs typically follows a ‘pre-training and fine-tuning’ paradigm [5]. However, both stages may inadvertently result in the model learning sensitive information or harmful content [6, 7, 8], which raises concerns about privacy violations, copyright infringement, or the generation of harmful content.
@@
To address the privacy or harmful content concerns, machine unlearning has been proposed as a method to remove the influence of specific data from a model without requiring full retraining [12]. Therefore, LLM unlearning has become a research hotspot [10, 11, 13]. It hypothesizes that the unlearned LLM is statistically indistinguishable from one retrained without the target data. Ideal LLM unlearning aims to identify and remove parameters associated with target knowledge, known as exact unlearning. However, this is limited by the poor interpretability of LLMs. Some efforts have shifted toward observable LLM outputs [10, 11, 14, 15], leading to a form of approximate unlearning. It assumes that the unlearned LLM exhibits different behavior from the original model on the target knowledge, while behaving similarly in the remaining knowledge. This is referred to as behavioral unlearning [16], where the goal is to hide target knowledge from the LLM’s behavior by modifying inputs or output logits.
@@
However, these LLM unlearning schemes have inherent weaknesses. First, the schemes involving fine-tuning LLM introduce considerable computational overhead, like gradient ascent, so they cannot adapt to frequent unlearning requests. Second, closed-source LLMs only provide their interfaces, making some schemes unsuitable. Third, LLMs possess emergent abilities, which means that forgetting must extend beyond specific data to include related content. Shumailov et al. [17] found that LLMs can recall previously forgotten knowledge through their in-context abilities, exposing a vulnerability in existing unlearning schemes to this ‘un-unlearning’ phenomenon. Furthermore, some unlearning schemes may reduce the overall utility of the model, potentially leading to catastrophic forgetting [18]. In Table I, we list three representative LLM unlearning schemes and their weaknesses.
@@
To tackle above weaknesses of LLM unlearning, we find that Retrieval-Augmented Generation (RAG) is a new emerged technology that does not need to change any parameter but with adjusting the output of LLMs [19, 20]. This special characteristic can be applied in machine unlearning. In general, the RAG workflow comprises two primary components: an information retrieval module and a text generation module [21, 22]. The retrieval module first retrieves relevant information from a knowledge base based on the input. The LLM then uses this retrieved information to generate the final response. This process can effectively improve the quality of outputs, reduce the likelihood of misinformation and hallucinations, and, most importantly, hide the target information to implement behavioral unlearning without revising the parameters.
@@
As illustrated in Figure 1 (a), RAG can revise the misinformation knowledge in the external knowledge base, effectively correcting it in LLM. When regarding misinformation as a forgotten target, the scalability of RAG technology offers a promising solution for LLM unlearning. In Figure 1 (b), through adding confidentiality requirements of the target knowledge in the knowledge base, RAG can hide the forgotten target’s information without requiring access to the unlearned LLM. Figure 1 (c) indicates that an ideal unlearning scheme would remove the target knowledge in LLM, thus not generating its related content. Fortunately, leveraging RAG to keep the forgotten target confidential serves a similar function to behavioral unlearning.
@@
Combined with the idea of modifying the input, we leverage RAG technology to propose a novel behavioral LLM unlearning framework. Specifically, it treats unlearning the target data as a constrained optimization problem for constructing its unlearned knowledge. The knowledge consists of a retrieval component and a constraint component. For each target, we use a heuristic method to build relevant knowledge (i.e., retrieval component) and confidentiality requirements (i.e., constraint component). To improve the effectiveness of unlearning, we apply prompt engineering to optimize both components. Our approach only modifies the external knowledge base to implement behavioral unlearning. This flexibility allows our framework to be applied to both open-source and closed-source models. Furthermore, we can utilize the target LLM to acquire additional knowledge relevant to the target data, thereby addressing the ‘un-unlearning’ phenomenon. Meanwhile, RAG serves as an an external plug-in that will not affect the model utility when inputs are not related to the knowledge base. This ensures that our scheme can easily land in various LLMs with minimal side effects and low overhead.
In our experiment, we focused on two key unlearning objectives: trained samples and learned concepts. RAG-based unlearning was implemented in both open-source and closed-source models, including GPT-4o, Gemini, and Llama-2-7b-chat. We evaluated our unlearning framework comprehensively across five key dimensions: effectiveness, universality, harmlessness, simplicity, and robustness, and found that the framework performed well in each dimension. Especially, we explored the resistance of our scheme against challenging and practical scenarios, such as jailbreak attacks and adaptive attacks. Furthermore, we extended RAG-based unlearning to multimodal large language models (MLLMs) and LLM-based agents, and achieved effective forgetting effects.
In summary, RAG-based unlearning enables model providers to control the knowledge boundaries of LLMs by managing external knowledge bases, thereby safeguarding privacy, protecting copyrighted data, and eliminating harmful content. Our primary contributions are as follows:
  • We are the first to propose a behavioral unlearning framework leveraging RAG technology, which can be applied to many LLMs.
@@ II-A1 The Definition of Machine Unlearning
There are two types of machine unlearning. The first is exact unlearning, and it guarantees that the distribution of $F_{U}$ and that of a model retrained on $D_{R}$ are indistinguishable. In practice, exact unlearning can only be implemented for tiny-scale models. Therefore, approximate unlearning comes up and only guarantees that the activation results of $F_{U}$ and that of a retrained model are indistinguishable. Clearly, LLMs are suitable for approximate unlearning due to their complex architectures. Notably, the output of LLMs differs from that of traditional models, Definition 1 cannot be extended to LLM unlearning. We will detail it in Section III-B.
@@ II-B1 Machine Unlearning
Machine unlearning has become a crucial technique for addressing various safety issues, such as privacy protection and copyright protection. It originates from the right to be forgotten, which grants users the right to request that a model provider remove the influence of its data from the model. Currently, machine unlearning has expanded into various fields, including image classification [24, 25], and graph neural networks [26]. In small-scale models, unlearning requests include sample unlearning [27], class unlearning [28], and feature unlearning [29]. To achieve these goals, data reorganization and model manipulation are employed. In the first approach, model providers will delete or modify the target data. For example, Bourtoule et al. developed a machine unlearning framework called SISA [24]. For model manipulation, model shifting, and model pruning are used to affect the knowledge boundary regarding the target data. For instance, Wang et al. [30] proposed a class-unlearning scheme by quantifying the class discrimination of different channels in convolutional neural networks.
@@ II-B2 Existing LLM Unlearning
Different from machine unlearning, LLM unlearning remains in an exploratory phase. There are three approximate LLM unlearning schemes: retraining-based, offset-based and input-based. In the first manner, Yao et al. [9] used a gradient ascent algorithm to increase the LLM’s loss on the target, causing the output to deviate from the original content. Similarly, Eldan et al. [13] re-labeled the forgotten data as non-sensitive responses, thereby minimizing their likelihoods. These unlearning schemes require accessing and manipulating LLMs and cannot be applied to closed-source LLMs. In the second manner, Huang et al. [10] employed logit shifting with two small-scale models to modify the LLM’s output. When given a forgotten sample, the shift prevents the LLM from generating sensitive information. However, this scheme relies on the performance of small models and requires precise logit shifting calculations. Input-based unlearning is popular, Pawelczyk et al. [11] leveraged the contextual capabilities of LLMs to implement an in-context unlearning method. They guided the model’s behavior by constructing counterfactual instances of the target data. However, this scheme increases the risk of privacy leaks, as it requires storing sensitive instances. Notably, the latter two unlearning schemes fall under behavioral unlearning, as they constrain the output of target knowledge without altering the LLM’s parameters.
@@ II-B3 Verification for Machine Unlearning
The purpose of verification is to ensure that the model cannot reveal information about the target data after unlearning operations. The model should closely resemble its counterpart that has never seen the target data. Currently, several verification methods for machine unlearning exist, including retraining time [24], privacy and security attacks [31], and accuracy [32]. The first manner measures how quickly the unlearned model can relearn the forgotten data. If the model can recover its original performance in a very short time, this suggests that it may still retain information about the forgotten data. Furthermore, several representative membership inference attacks (MIAs) were applied to verify machine unlearning, such as LOSS [33], ZLib [34], GradNorm [35] and MinK [36]. In a successful demonstration of unlearning, MIAs determine the target data to be ‘in’ for the pre-unlearned model and ‘out’ for the post-unlearned model. Similarly, Guo et al. [31] used backdoored data as the target and calculated the drop of attack success rate. Literature [32] used an accuracy-based verification method, comparing the output of the model after unlearning with that of a model that has never seen the target data. Prior studies [11, 35] have used MIAs to assess residual memorization in LLMs. We also applied this method to evaluate the effectiveness of LLM unlearning, as detailed in Section V-A.
@@ II-B4 Retrieval-Augmented Generation
RAG in LLMs is an enhancement method that combines retrieval and generation techniques. Compared to regular LLMs, RAG-based models are effective at handling complex tasks, especially in contexts that require detailed background knowledge and specific facts. Specifically, by introducing a retrieval process, RAG-based models can incorporate the most relevant information related to input prompts, reducing factual errors during the generation process. Additionally, RAG techniques are well-suited for handling multi-turn dialogues [37]. Currently, some researchers improved RAG techniques regarding retrieval process and generation process. Jiang et al. [38] optimized the retrieval algorithm for faster and more accurate information extraction. While Siriwardhana et al. [39] focused on the generation process to better integrate and utilize the retrieved knowledge.
@@ III Problem Definition
BKBK The benign knowledge base
UKUK The unlearned knowledge base
@@ III-A Notations
III-B The Definition of Behavioral LLM Unlearning
Due to the poor interpretability of LLMs, relying on internal parameters or activation values for unlearning is impractical, such as exact unlearning [24]. Following recent studies [10, 11], we focused instead on the model’s observable output behavior, known as behavioral unlearning. This assumes that the unlearned LLM behaves differently from the original model on target knowledge, while remaining consistent on unrelated knowledge. LLM unlearning involves two primary objectives: sample-level [11] and concept-level [9] knowledge. We define behavioral unlearning with respect to each.
@@ III-B The Definition of Behavioral LLM Unlearning
Notably, behavioral unlearning is not intended to replace other unlearning schemes. Rather, it can serve as a component in privacy protection and content copyright control, complementing other schemes. In this work, we adopt RAG to constrain the behavior of the LLM, achieving approximate LLM unlearning.
@@ III-C Threat Model
In contrast to traditional unlearning, we consider real-world unlearning scenarios for LLMs, including privacy protection, copyright compliance, and harmful content removal. The first scenario primarily involves sample unlearning, whereas the latter two involve concept unlearning. As illustrated in Figure 3, both types of requests typically arise after model deployment. In such cases, a data contributor may request the withdrawal of specific samples (partial training data) or concepts (sets of related training data). The model provider then applies behavioral unlearning schemes to suppress the model’s ability to generate responses containing the target knowledge. After unlearning, normal users can no longer access this knowledge through the LLM’s outputs. The unlearned model must be verified to ensure the target information has been hidden. Under this threat model, we detail the capabilities of the user and the model provider.
  • Model Provider. For sample unlearning, the provider can identify the specific samples to be forgotten, but does not require access to the LLM’s parameters. For concept unlearning, the provider knows only the target concepts, without needing the corresponding samples. In both cases, the provider can integrate an external RAG system into the LLM to augment the input prompts and faithfully follow the RAG pipeline. In this context, the provider has full control over the retriever configuration, knowledge base construction, and prompt template design. This assumption supports the applicability of our scheme to a wide range of real-world LLMs.
  • Normal Users. They interact with the LLM through its interface, submitting arbitrary text inputs to make the LLM respond as instructed. Moreover, they cannot access LLM parameters, intermediate activations, or bypass the RAG pipeline (e.g., direct access to the base LLM is disallowed). Under this setting, the behavioral suppression we adopt completely blocks access to target knowledge, ensuring valid unlearning.